# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 8 Acknowledgements
$\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}-\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k+1}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k},$ (23)
where the single-valuedness of the right-hand side has been shown in the proof of Lemma 2.10. Let $q_{\mathsf{r}}^{k}:=\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}$, $\mathcal{B}:=\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}$, then, (23) is rewritten as (by maximal monotonicity of $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ and celebrated Moreau’s identity):
@@
It is known from the proof of Theorem 5.3 that $q_{\mathsf{r}}^{k}-q_{\mathsf{r}}^{k+1}\rightarrow 0$ and $q_{\mathsf{r}}^{k}\rightharpoonup q^{\star}$, as $k\rightarrow\infty$. Since $\mathcal{B}^{-1}$ is also maximally monotone (by [4, Proposition 20.22]), then $\mathsf{gra}\mathcal{B}^{-1}$ is sequentially closed in $(\mathsf{ran}\mathcal{Q})^{\text{weak}}\times(\mathsf{ran}\mathcal{Q})^{\text{strong}}$, i.e., $(\mathsf{ran}\mathcal{Q},\|\cdot\|_{\mathcal{Q}})\times(\mathsf{ran}\mathcal{Q},\langle\cdot|\cdot\rangle_{\mathcal{Q}})$, by [4, Proposition 20.38]. Thus, we obtain
$0\in\mathcal{B}^{-1}q^{\star}\Longrightarrow q^{\star}\in\mathcal{B}(0)=\sqrt{\mathcal{Q}}\mathcal{A}^{-1}(0)\Longrightarrow x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}(0)=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A}).$
@@
Under Assumptions 1 and 2, given $\mathcal{T}$ in (2), then $\forall x\in\mathsf{ran}\mathcal{Q}$, $\exists\lambda_{\inf},\lambda_{\sup}>0$, such that
@@
(iii) $\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x_{1}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x_{2}\|\leq\sqrt{\frac{\lambda_{\sup}}{\lambda_{\inf}}}\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$, $\forall x_{1},x_{2}\in\mathsf{dom}\mathcal{T}$.
@@
(i)$\Longrightarrow$(iii):
$\|(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x_{1}-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x_{2}\|$
$=$ $\|y_{1}-y_{2}\|=\|y_{1,\mathsf{r}}+y_{1,\mathsf{k}}-y_{2,\mathsf{r}}-y_{2,\mathsf{k}}\|$
$\leq$ $\|y_{1,\mathsf{r}}-y_{2,\mathsf{r}}\|+\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|$
[…]
(iii)$\Longrightarrow$(i):
$\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|$$\leq$ $\|y_{1}-y_{2}\|+\|y_{1,\mathsf{r}}-y_{2,\mathsf{r}}\|$
[…]

### v4

@@ 8 Acknowledgements
$\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}-\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k+1}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathrm{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k},$ (22)
where the single-valuedness of the right-hand side has been shown in the proof of Lemma 2.11. Let $q_{\mathsf{r}}^{k}:=\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}$, $\mathcal{B}:=\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathrm{ran}\mathcal{Q}}$, then, (22) is rewritten as (by maximal monotonicity of $\mathcal{B}:\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$ and celebrated Moreau’s identity):
@@
It is known from the proof of Theorem 5.3 that $q_{\mathsf{r}}^{k}-q_{\mathsf{r}}^{k+1}\rightarrow 0$ and $q_{\mathsf{r}}^{k}\rightharpoonup q^{\star}$, as $k\rightarrow\infty$. Since $\mathcal{B}^{-1}$ is also maximally monotone (by [5, Proposition 20.22]), then $\mathrm{gra}\mathcal{B}^{-1}$ is sequentially closed in $(\mathrm{ran}\mathcal{Q})^{\text{weak}}\times(\mathrm{ran}\mathcal{Q})^{\text{strong}}$, i.e., $(\mathrm{ran}\mathcal{Q},\|\cdot\|_{\mathcal{Q}})\times(\mathrm{ran}\mathcal{Q},\langle\cdot|\cdot\rangle_{\mathcal{Q}})$, by [5, Proposition 20.38]. Thus, we obtain
$0\in\mathcal{B}^{-1}q^{\star}\Longrightarrow q^{\star}\in\mathcal{B}(0)=\sqrt{\mathcal{Q}}\mathcal{A}^{-1}(0)\Longrightarrow x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{A}^{-1}(0)=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}(\mathrm{zer}\mathcal{A}).$
@@
Under Assumptions 1 and 2, given $\mathcal{T}$ in (2), then $\forall x\in\mathrm{ran}\mathcal{Q}$, $\exists\lambda_{\inf},\lambda_{\sup}>0$, such that
@@
(iii) $\|\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{T}x_{1}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\mathcal{T}x_{2}\|\leq\sqrt{\frac{\lambda_{\sup}}{\lambda_{\inf}}}\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$, $\forall x_{1},x_{2}\in\mathrm{dom}\mathcal{T}$.
@@
The equivalence bewtween (i) and (iii) immediately follows from Fact 4.